# How Effective Are Time-Series Models for Rainfall Nowcasting? A Comprehensive Benchmark for Rainfall Nowcasting Incorporating PWV Data

### v2

@@ Introduction
Specifically, for the JFNG station, we utilize a real-world meteorological dataset spanning from February 14, 2018, 00:00 to December 31, 2022, 23:00, with observations recorded every 15 minutes, totaling 171,070 time steps without missing entries. Each record consists of six variables (excluding the timestamp): five meteorological features and one target variable representing rainfall. Figure 2 illustrates the distribution of values for each variable. Specifically, the input features include:
@@
  • tp: total precipitation (target), obtained from ERA5 reanalysis data.
Correlation Analysis
@@
Across all correlation matrices, PWV shows the strongest positive correlation with tp (e.g., a Pearson coefficient of 0.27), highlighting its value as an informative feature for short-term rainfall prediction. In contrast, variables like t2m and sp exhibit weak or negative correlations, indicating limited relevance at nowcasting timescales. Since rainfall nowcasting focuses on the next 0 to 3 hours, model performance hinges on features that reflect rapid and physically meaningful atmospheric changes. Among them, PWV emerges as the most reliable indicator of imminent rainfall, making its inclusion essential in both data selection and model design. Moreover, numerous meteorological studies have observed that once PWV exceeds a certain threshold, the probability of precipitation increases significantly (Deng et al. 2021; Papée and Montefinale 1961).
Figure 4: Analysis of Data Properties in RainfallBench
Analysis of Data Properties
@@
Zero Inflation: The majority of target values are zeros, reflecting the sparse and event-driven nature of rainfall. This sparsity undermines conventional modeling assumptions that rely on frequent signal continuity. Figure 4(a) shows the distribution of 500 randomly sampled records. It is evident that the majority have a tp value of 0. In total, there are 121,572 such records, accounting for 71% of the entire dataset.
@@
Comparison Baselines
To ensure a comprehensive benchmark evaluation, we selected 20 models spanning commonly used architectures, including MLP-based, CNN/TCN-based, RNN-based, GNN-based, KAN-based, and Transformer-based designs. To maintain both relevance and rigor, all selected models are state-of-the-art methods proposed in top-tier AI conferences within the past four years. Details of the selected models are summarized in Table 1.
Evaluation Strategy
@@
$\mathcal{L}_{\text{out}}=\{4,6,8,10,12\}$
corresponding to 1 to 3 hour forecasts in the nowcasting task (15-minute resolution). Each model is evaluated on all combinations from the Cartesian product $\mathcal{L}_{\text{in}}\times\mathcal{L}_{\text{out}}$.
@@
Accuracy and Reliability Evaluation. Two separate leaderboards are reported accordingly: one for average MSE-based ranking and the other for average MAE-based ranking across all input-output configurations, providing a comprehensive view of each model’s performance in terms of both accuracy and error scale sensitivity.
Extreme Rainfall Evaluation. Accurate extreme rainfall forecasting is vital for disaster mitigation. In RainfallBench, we follow the T/CMSA 0013-2019 standard  [^2] https://www.chinamsa.org/, under which extreme rainfall is defined as any 15-minute period with precipitation exceeding 2mm. We label these intervals and evaluate models using MSE and MAE computed only on them.
$\text{MSE}_{\text{extreme}}$ $=\frac{1}{|E|}\sum_{t\in E}(\hat{y}_{t}-y_{t})^{2}$ (5)
$\text{MAE}_{\text{extreme}}$ $=\frac{1}{|E|}\sum_{t\in E}|\hat{y}_{t}-y_{t}|$ (6)
where $E$ denotes the set of time steps labeled as extreme rainfall.
Bi-Focus Precipitation Forecaster
Motivation
Our benchmark shows that while existing time series models often consider non-stationarity, they struggle with rainfall nowcasting due to two overlooked domain-specific challenges: zero inflation and temporal decay.
@@
Figure 5: Bi-Focus Precipitation Forecaster
Non-Zero Focus
@@
Table 1: Comparison of state-of-the-art methods. The red indicates the best-performing model, while the pink highlights the second-best.
Methods Publication 24(4) 24(6) 24(8) 24(10) 24(12) 48(4) 48(6) 48(8) 48(10) 48(12)
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
MLP-based
DLinear (Zeng et al. 2023) AAAI 2023 0.0099 0.0247 0.0189 0.0361 0.0282 0.0463 0.0371 0.0561 0.0465 0.0658 0.0099 0.0243 0.0188 0.0358 0.0281 0.0461 0.0370 0.0550 0.0450 0.0631
Koopa (Liu et al. 2023b) NIPS 2023 0.0447 0.1859 0.0552 0.1928 0.0703 0.2012 0.0804 0.2054 0.0893 0.2102 0.0439 0.1833 0.0530 0.1884 0.0633 0.1946 0.0727 0.2006 0.0822 0.2045
TimeMixer (Wang et al. 2024a) ICLR 2024 0.1616 0.2332 0.1096 0.2214 0.0989 0.2168 0.3193 0.2654 0.1884 0.2633 0.0392 0.1770 0.0627 0.1985 0.6356 0.3574 0.0721 0.1996 0.0796 0.2019
FilterTS (Wang et al. 2025b) AAAI 2025 0.0215 0.0738 0.0323 0.0841 0.0546 0.1738 0.0517 0.0907 0.0640 0.1266 0.0233 0.1018 0.0404 0.1555 0.0417 0.0983 0.0519 0.1019 0.0736 0.1889
RNN-based
SegRNN (Lin et al. 2023) Arxiv 2023 0.0098 0.0239 0.0201 0.0390 0.0277 0.0440 0.0379 0.0502 0.0453 0.0548 0.0086 0.0229 0.0168 0.0323 0.0252 0.0428 0.0335 0.0449 0.0435 0.0592
xLSTM (Beck et al. 2025) NIPS 2024 0.0104 0.0318 0.0183 0.0474 0.0280 0.0491 0.0360 0.0596 0.0451 0.0662 0.0096 0.0293 0.0191 0.0442 0.0290 0.0503 0.0362 0.0635 0.0466 0.0692
P-sLSTM (Kong et al. 2025) AAAI 2025 0.0083 0.0228 0.0163 0.0325 0.0242 0.0369 0.0328 0.0469 0.0401 0.0515 0.0081 0.0198 0.0163 0.0296 0.0262 0.0526 0.0333 0.0454 0.0413 0.0532
TCN&CNN-based
TimesNet (Wu et al. 2023) ICLR 2023 0.0482 0.1873 0.0560 0.1938 0.0720 0.2008 0.0845 0.2071 0.0953 0.2113 0.0482 0.1839 0.0618 0.1924 0.0753 0.1998 0.0892 0.2047 0.0981 0.2096
TimeMixer++ (Wang et al. 2025a) ICLR 2025 0.0406 0.1747 0.0558 0.1879 0.0644 0.1925 0.0773 0.1995 0.0875 0.2058 0.0416 0.1748 0.0513 0.1824 0.0651 0.1917 0.0751 0.1973 0.0850 0.2041
xPatch (Stitsyuk and Choi 2025) AAAI 2025 0.0413 0.1806 0.0487 0.1758 0.0587 0.1804 0.0718 0.1982 0.0795 0.1828 0.0266 0.1307 0.0512 0.1870 0.0603 0.1922 0.0691 0.1959 0.0612 0.1050
GNN-based
MSGNet (Cai et al. 2024) AAAI 2024 0.0484 0.1867 0.0607 0.1939 0.0726 0.1998 0.0840 0.2059 0.0965 0.2107 0.0477 0.1825 0.0599 0.1898 0.0720 0.1961 0.0875 0.2027 0.0967 0.2070
TimeFilter (Hu et al. 2025b) ICML 2025 0.0409 0.1812 0.0509 0.1884 0.0630 0.1942 0.0727 0.2004 0.0841 0.2058 0.0383 0.1757 0.0480 0.1827 0.0570 0.1890 0.0670 0.1951 0.0753 0.1991
KAN-based
TimeKAN (Huang et al. 2025) ICLR 2025 0.0425 0.1834 0.0527 0.1912 0.0649 0.1979 0.0751 0.2032 0.0851 0.2074 0.0406 0.1801 0.0499 0.1857 0.0599 0.1924 0.0698 0.1981 0.0806 0.2038
MMK (Han et al. 2024) Arxiv 2024 0.0275 0.0178 0.0407 0.0361 0.0511 0.0553 0.0612 0.0765 0.0694 0.0959 0.0307 0.0197 0.0480 0.0433 0.0590 0.0667 0.0612 0.0727 0.0748 0.1020
Transformer-based
Informer (Zhou et al. 2021) AAAI 2021 0.0093 0.0281 0.0176 0.0365 0.0298 0.0471 0.0358 0.0536 0.0555 0.0738 0.0102 0.0262 0.0191 0.0399 0.0270 0.0440 0.0347 0.0522 0.0462 0.0633
PatchTST (Nie et al. 2023) ICLR 2023 0.0429 0.1823 0.0540 0.1907 0.0688 0.1997 0.0780 0.2039 0.0886 0.2081 0.0411 0.1774 0.0521 0.1874 0.0608 0.1925 0.0724 0.1955 0.0819 0.2011
iTransformer (Liu et al. 2024) ICLR 2024 0.0411 0.1752 0.0512 0.1832 0.0647 0.1906 0.0758 0.1973 0.0876 0.2028 0.0395 0.1729 0.0505 0.1804 0.0641 0.1920 0.0742 0.1966 0.0886 0.2038
TimeXer (Wang et al. 2024b) NIPS 2024 0.0584 0.1879 0.0719 0.1969 0.0862 0.2042 0.0950 0.2085 0.1043 0.2149 0.0413 0.1780 0.0924 0.1984 0.0617 0.1913 0.0736 0.1975 0.0801 0.2017
PPDformer (Wan et al. 2025) ICASSP 2025 0.0222 0.0790 0.0325 0.0855 0.0449 0.0927 0.0558 0.0990 0.0668 0.1067 0.0246 0.0848 0.0492 0.1765 0.1765 0.1000 0.0803 0.2011 0.0850 0.2051
Informer(with BFPF) Ours 0.0104 0.0276 0.0176 0.0323 0.0263 0.0428 0.0392 0.0525 0.0428 0.0531 0.0093 0.0231 0.0172 0.0355 0.0263 0.0404 0.0341 0.0500 0.0425 0.0593
Temporal Focus
@@
Experiments
Experiment Setup
@@
All experiments are conducted on an NVIDIA H100 80GB GPU. To ensure evaluation consistency, all results are computed using de-normalized actual rainfall values. For robustness, each experiment is repeated three times, and the average performance is reported as the final result. The datasets were split into training, validation, and test sets in a 7:1:2 ratio.
Experiment Results
Comprehensive forecasting results are listed in Table 1 with the best in red and the second in pink. The lower MSE/MAE indicates the more accurate prediction result. Surprisingly, RNN-based models achieve consistently strong results, with P-sLSTM demonstrating the best overall performance across all experimental settings. Informer with our BFPF outperforms all Transformer-based counterparts, achieving the best performance under the extreme rainfall setting among all baselines.
@@
RQ1: How do different model architectures perform in rainfall nowcasting, and which type achieves the best results? (Forecast Reliability and Accuracy Evaluation)
RQ2: How does model performance vary with changes in forecasting horizon and predict length? (Multi-temporal scale Evaluation)
RQ3: How do different models perform for extreme rainfall forecasting? (Extreme Rainfall Evaluation)
RQ4: Does Incorporating PWV Significantly Enhance Rainfall Nowcasting?
@@
RQ5: Do our proposed Bi-Focus Precipitation Forecaster improve the performance of Transformer-based models?
RQ1: Forecast Reliability and Accuracy Evaluation
As shown in Figure 6, two of the top three models are RNN-based, suggesting that RNNs are particularly well-suited for rainfall nowcasting. This is largely due to the nature of our RainfallBench dataset, which features high zero-inflation and rapid temporal decay—properties that favor models with strong short-term memory. RNNs, with their ability to emphasize recent inputs and handle sparse, irregular sequences, align well with these data characteristics. This explains why models like P-sLSTM consistently outperform others in our benchmark.
RQ2: Multi-temporal scale Evaluation
Figure 7: Model Performance Comparison Across Multiple Temporal Scales (MSE).
To ensure a fair and comprehensive evaluation, we selected representative models from six widely-used architectural families in time-series forecasting.
We consider two evaluation settings. In the first setting, we fix the input length to 24 steps and vary the forecast horizon. As expected, model errors increase with longer prediction lengths. Among all models, P-sLSTM achieves the best performance, closely followed by our model. Both significantly outperform baseline models, whose errors grow more rapidly.
In the second setting, we fix the output length to 12 steps and vary the input length (24 vs. 48). We observe that, for most models, increasing the input horizon leads to lower MSE, indicating that a longer historical context improves forecasting performance.
RQ3: Extreme Rainfall Evaluation
Figure 8: Model Performance on Extreme Rainfall Prediction.
As shown in Figure 8, our proposed model achieves the best performance in extreme rainfall prediction, attributed to the BFPF’s enhanced capability in capturing sparse signals. Interestingly, compared with general settings (Figure 6), we observe consistently higher MSE under extreme rainfall conditions, revealing the limitations of existing models and the need for advances in extreme rainfall forecasting. Comprehensive results can be found in Supplementary Material.
RQ4: Effect Analysis of PWV
To verify the effectiveness of the PWV variable in rainfall nowcasting, we train our proposed model using datasets with (w/) and without (w/o) PWV. As shown in Table 2, incorporating PWV significantly reduces nowcasting errors, confirming both its importance for the task and the necessity of including it in RainfallBench.
Table 2: Ablation analysis of PWV
Module 48(4) 48(6) 48(8) 48(10) 48(12)
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Ours(w/o PWV) 0.0094 0.0281 0.0194 0.0370 0.0272 0.0461 0.0345 0.0462 0.0428 0.0530
Ours(w/ PWV) 0.0093 0.0231 0.0172 0.0355 0.0263 0.0404 0.0341 0.0500 0.0425 0.0593
RQ5: Effect Analysis of Bi-Focus Precipitation Forecaster

### v3

@@ III-D1 Data Overview
Specifically, for the HKST station, we utilize a real-world meteorological dataset spanning from January 1, 2018, 00:00 to January 1, 2024, 00:00, with observations recorded every 1 hour, totaling 52,585 time steps without missing entries. Each record consists of six variables (excluding the timestamp): five meteorological features and one target variable representing rainfall. Figure 3 illustrates the distribution of values for each variable. Specifically, the input features include:
@@
  • tp: total precipitation (target), obtained from GPM IMERG.
III-D2 Correlation Analysis
@@
Across all correlation matrices, PWV shows the strongest positive correlation with tp (e.g., a Pearson coefficient of 0.27), highlighting its value as an informative feature for short-term rainfall prediction. In contrast, variables like t2m and sp exhibit weak or negative correlations, indicating limited relevance at nowcasting timescales. Since precipitation nowcasting focuses on the next 0 to 6 hours, model performance hinges on features that reflect rapid and physically meaningful atmospheric changes. Among them, PWV emerges as the most reliable indicator of imminent rainfall, making its inclusion essential in both data selection and model design. Moreover, numerous meteorological studies have observed that once PWV exceeds a certain threshold, the probability of precipitation increases significantly [28, 29].
Figure 5: Analysis of data properties in RainfallBench. The benchmark exhibits three key characteristics distinguishing it from standard time series: (i) zero inflation, (ii) temporal dependency decay, and (iii) non-stationarity, along with their implications for modeling.
III-D3 Analysis of Data Properties
@@
Zero Inflation: The majority of target values are zeros, reflecting the sparse and event-driven nature of rainfall. This sparsity undermines conventional modeling assumptions that rely on frequent signal continuity. Figure 5(a) shows the distribution of 500 randomly sampled records. It is evident that the majority have a tp value of 0. In total, there are 43,313 such records, accounting for 82.3% of the entire dataset.
@@
III-E Comparison Baselines
To ensure a comprehensive benchmark evaluation, we selected 17 models spanning commonly used architectures, including MLP-based, CNN/TCN-based, RNN-based, GNN-based, KAN-based, and Transformer-based designs. To maintain both relevance and rigor, all selected models are state-of-the-art methods proposed in top-tier AI conferences within the past four years. Details of the selected models are summarized in Table II.
III-F Evaluation Strategy
@@
$\mathcal{L}_{\text{out}}=\{2,4,6\}$
corresponding to 1 to 6 hour forecasts in the nowcasting task (1-hour resolution). Each model is evaluated on all combinations from the Cartesian product $\mathcal{L}_{\text{in}}\times\mathcal{L}_{\text{out}}$. The forecast sequence length $L_{\text{out}}$ is defined as an element of the output length set: $L_{\text{out}}\in\mathcal{L}_{\text{out}}.$
@@
Multi-Forecast Resolution Evaluation. In this evaluation, we keep the input data at a fixed temporal resolution of 1 hour, while assessing model performance under different forecast resolutions. Formally, we define the set of forecast resolutions as
$\mathcal{R}_{\text{out}}=\{1\text{h},2\text{h},3\text{h}\}.$
$L_{\text{out}}=\frac{H}{\mathcal{R}_{\text{out}}}$
Each model is evaluated at each forecast resolution in the set $\mathcal{R}_{\text{out}}$.
For each forecast resolution, we compute both the MSE and MAE between the predicted rainfall sequence $\hat{y}_{1:L_{\text{out}}}$ and the ground truth sequence $y_{1:L_{\text{out}}}$, defined as:
$\text{MSE}=\frac{1}{L_{\text{out}}}\sum_{t=1}^{L_{\text{out}}}(\hat{y}_{t}-y_{t})^{2},$ (5)
$\text{MAE}=\frac{1}{L_{\text{out}}}\sum_{t=1}^{L_{\text{out}}}|\hat{y}_{t}-y_{t}|.$ (6)
This evaluation allows us to analyze the model’s robustness across different forecast time granularities and to understand how the choice of output temporal resolution affects rainfall prediction performance.
Extreme Rainfall Evaluation.
Accurate extreme rainfall forecasting is vital for disaster mitigation. In RainfallBench, we follow the T/CMSA 0013-2019 standard [^2] http://www.chinamsa.org/uploads/file/20191106142922_61962.pdf, under which extreme rainfall is defined as any hourly period with precipitation exceeding 4 mm. We label these intervals and evaluate model performance using the Extreme Event Reconstruction Error (EERE) and its absolute variant (AEERE), computed only over extreme rainfall periods:
$\text{EERE}=\frac{1}{|E|}\sum_{t\in E}(\hat{y}_{t}-y_{t})^{2}$ (7)
$\text{AEERE}=\frac{1}{|E|}\sum_{t\in E}|\hat{y}_{t}-y_{t}|$ (8)
where $E$ denotes the set of time steps labeled as extreme rainfall events. A lower EERE or AEERE indicates better reconstruction fidelity of high-intensity precipitation patterns.
IV Bi-Focus Precipitation Forecaster
IV-A Motivation
Our benchmark shows that while existing time series models often consider non-stationarity, they struggle with precipitation nowcasting due to two overlooked domain-specific challenges: zero inflation and temporal decay.
@@ IV-A Motivation
Figure 6: Overview of the BFPF module for transformer-based rainfall forecasting. It consists of two key components: (i) Non-Zero Focus, which reduces distractions from non-rainy periods, and (ii) Temporal Focus, which emphasizes temporally proximate context to improve prediction accuracy.
IV-B Non-Zero Focus
@@
IV-C Temporal Focus
@@
TABLE II: Comparison of state-of-the-art methods. The red indicates the best-performing model, while the pink highlights the second-best. Results are obtained with an input sequence length of 24 and an output sequence length of 6.
Methods Publication J340 ZIMM P095 MTLA ARTA BFTA FLM5 Average
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
MLP-based
DLinear AAAI 2023 1.3654 0.4574 2.0124 0.4471 0.3804 0.1159 1.6673 0.4204 4.0069 0.5785 0.8255 0.2680 0.0000 0.0000 1.4654 0.3250
Koopa NIPS 2023 1.4559 0.3525 2.0211 0.3716 0.3896 0.1030 1.7539 0.3271 4.0590 0.4896 0.8442 0.2272 0.0000 0.0000 1.5034 0.2673
FilterTS AAAI 2025 1.4796 0.3573 2.0549 0.3656 0.3905 0.1064 1.7334 0.3224 4.1174 0.4916 0.8438 0.2242 0.0000 0.0000 1.5171 0.2668
RNN-based
SegRNN Arxiv 2023 1.3102 0.4228 2.0623 0.4326 0.4810 0.1102 1.6353 0.3850 4.3240 0.5598 0.8413 0.2634 0.0000 0.0000 1.5220 0.3105
xLSTM NIPS 2024 1.3002 0.4240 1.9704 0.3838 0.3625 0.1144 1.5989 0.4228 3.7109 0.5639 0.7573 0.2500 0.0000 0.0044 1.3857 0.3090
P-sLSTM AAAI 2025 1.3368 0.4073 2.0091 0.4278 0.3693 0.1096 1.6303 0.3777 3.8603 0.5651 0.7979 0.2584 0.0000 0.0000 1.4277 0.3066
TCN&CNN-based
TimesNet ICLR 2023 1.4781 0.3538 2.0045 0.3700 0.3930 0.1025 1.7046 0.3336 4.0875 0.4962 0.8470 0.2329 0.0000 0.0000 1.5021 0.2700
TimeMixer++ ICLR 2025 1.4825 0.3416 2.0404 0.3560 0.3906 0.0956 1.7141 0.3243 4.0926 0.4754 0.8567 0.2245 0.0000 0.0000 1.5110 0.2596
xPatch AAAI 2025 1.4169 0.3261 2.0269 0.3456 0.3815 0.0924 1.6956 0.3245 4.0443 0.4647 0.8354 0.2143 0.0000 0.0000 1.4858 0.2525
GNN-based
MSGNet AAAI 2024 1.4417 0.3451 1.9761 0.3657 0.3828 0.1042 1.6934 0.3279 4.0096 0.4854 0.8436 0.2262 0.0000 0.0000 1.4782 0.2650
TimeFilter ICML 2025 1.4128 0.3247 2.0217 0.3510 0.3857 0.0936 1.6920 0.3078 4.0255 0.4585 0.8318 0.2141 0.0000 0.0000 1.4814 0.2499
KAN-based
TimeKAN ICLR 2025 1.4101 0.3217 2.0198 0.3495 0.3780 0.0933 1.7112 0.3179 4.0501 0.4641 0.8361 0.2141 0.0000 0.0000 1.4865 0.2515
Transformer-based
Informer AAAI 2021 1.3601 0.3707 2.0258 0.5461 0.3913 0.1287 1.5952 0.3995 3.8286 0.6883 0.7757 0.2470 0.0000 0.0006 1.4252 0.3401
PatchTST ICLR 2023 1.4298 0.3230 2.0281 0.3538 0.3813 0.0926 1.7120 0.3105 4.0303 0.4582 0.8363 0.2212 0.0000 0.0000 1.4883 0.2513
iTransformer ICLR 2024 1.4752 0.3462 2.0870 0.3705 0.3909 0.1000 1.7288 0.3244 4.0953 0.4811 0.8759 0.2497 0.0000 0.0000 1.5217 0.2674
TimeXer NIPS 2024 1.6521 0.3806 2.1916 0.3848 0.3860 0.1025 1.8817 0.3563 4.2324 0.5082 0.8942 0.2417 0.0000 0.0000 1.6054 0.2819
PPDformer ICASSP 2025 1.4316 0.3390 2.0907 0.3666 0.3900 0.0971 1.7307 0.3216 4.1038 0.4734 0.8451 0.2354 0.0000 0.0000 1.5131 0.2618
Informer(with BFPF) Ours 1.3671 0.3878 2.2093 0.4783 0.0901 0.0565 1.2203 0.3324 3.3670 0.6555 0.4860 0.1929 0.0000 0.0003 1.2485 0.2730
Average \ 1.4226 0.3656 2.0473 0.3926 0.3730 0.1010 1.6722 0.3464 4.0025 0.5199 0.8152 0.2336 0.0000 0.0003 \ \
V Experiments
V-A Experiment Setup
@@
All experiments are conducted on an NVIDIA H100 80GB GPU. To ensure the generalizability of the experiment, we selected a representative station from each continent for the study. To ensure evaluation consistency, all results are computed using de-normalized actual rainfall values. For robustness, each experiment is repeated three times, and the average performance is reported as the final result. The datasets were split into training, validation, and test sets in a 7:1:2 ratio.
The seven GNSS stations used in experiments are geographically distributed across different continents, elevations, and climate zones. Their details are summarized below:
J340 (34.406°N, 135.364°E, 91.983 m) – Located in the Kinki region of Japan, this station is situated at a low elevation. It is characterized by a humid subtropical climate (Cfa), with four distinct seasons, hot and rainy summers, mild winters, and relatively evenly distributed precipitation throughout the year. The region is also occasionally affected by typhoons in summer. This station has a completeness rate of 99.82%.
ZIMM (46.877°N, 7.465°E, 956.34 m) – Located in central Switzerland in the Alps, this mid-altitude station exhibits a temperate continental climate (Dfb), characterized by cold, snowy winters and warm, humid summers, with precipitation distributed throughout the year and frequent summer thunderstorms. This station has a completeness rate of 99.85%.
P095 (39.698°N, -119.537°W, 1608.804 m) – Located in western Nevada, USA, this high-altitude station experiences a temperate desert climate (BWk), with arid conditions, large diurnal temperature variations, cold winters, hot summers, and low annual precipitation. This station has a completeness rate of 99.82%.
MTLA (-15.228°S, -59.35°W, 267.63 m) – Located in Mato Grosso, Brazil, this low-elevation station exhibits a tropical wet and dry climate (Aw/Am), with a pronounced wet season in summer (November–March) and a dry winter season. The region experiences high average annual temperatures. This station has a completeness rate of 99.64%.
ARTA (-38.618°S, 176.136°E, 369.779 m) – This station in the eastern North Island of New Zealand lies at a moderate elevation. The climate is temperate oceanic (Cfb), with mild and humid conditions year-round, evenly distributed precipitation, warm summers, and cool winters, strongly influenced by the surrounding ocean. This station has a completeness rate of 99.48%.
BFTA (-29.111°S, 26.205°E, 1441.266 m) – Situated in northern South Africa, this high-altitude station experiences a subtropical highland climate (Cwb), with warm and wet summers, cool and dry winters, and most precipitation occurring during the summer months. This station has a completeness rate of 98.16%.
FLM5 (-77.533°S, 160.271°E, 1869.726 m) – Situated on Mount Fleming, Antarctica, this high-elevation polar station is characterized by an ice cap climate (EF). It experiences extremely cold temperatures year-round, very low precipitation (mostly snow), strong winds, and a permanently frozen environment. This station has a completeness rate of 99.77%.
V-B Experiment Results
This analysis is based on the performance of various time series forecasting models in predicting rainfall at multiple GNSS stations (J340, ZIMM, P095, MTLA, ARTA, BFTA, FLM5). The evaluation metrics include MSE and MAE, where smaller values indicate higher prediction accuracy of the models. Comprehensive forecasting results are listed in Table II with the best in red and the second in pink. The lower MSE/MAE indicates the more accurate prediction result.
@@ V-B Experiment Results
RQ1: How do different model architectures perform in precipitation nowcasting, and which type achieves the best results?(Analysis from the Model Perspective)
RQ2: How dose various model performance in each area? (Analysis from the Dataset Perspective)
@@
RQ3: Do our proposed Bi-Focus Precipitation Forecaster improve the performance of Transformer-based models? (Effect Analysis of Bi-Focus Precipitation Forecaster)
RQ4: How does model performance vary with changes in forecasting horizon and predict length? (Multi-temporal scale Evaluation)
RQ5: How does model performance vary with changes in multi-forecast resolution? (Multi-Forecast Resolution Evaluation)
RQ6: How do different models perform for extreme rainfall forecasting? (Extreme Rainfall Evaluation)
V-C RQ1: Analysis from the Model Perspective
The MLP-based models, such as DLinear, Koopa, and FilterTS, exhibit moderate performance across most sites, particularly at the J340 and P095 stations, where the MSE and MAE values are relatively high, indicating noticeable prediction errors. RNN-based models, including SegRNN, xLSTM, and P-slSTM, demonstrate superior performance at several sites, particularly at the J340 and P095 stations. Among these, the xLSTM model achieves the lowest MSE, suggesting its high prediction accuracy at six stations. TCN and CNN-based models, such as TimesNet, TimeMixer+, and xPatch, perform exceptionally well at the ZIMM and P095 station, with the xPatch model yielding the lowest MAE values, highlighting its superior performance at this site. GNN-based models, including MSGNet and TimeFilter, show good performance at the MTLA and BFTA stations, with the TimeFilter model achieving the lowest MAE values at the ZIMM station. The KAN-based model, TimeKAN, demonstrates outstanding performance across multiple sites, especially at the J340, BFTA and P095 stations, where it records the lowest MAE values, indicating the highest prediction accuracy at these sites. Transformer-based models also exhibit strong performance at multiple sites, with the Informer model achieving the second lowest average MSE among all the models. PatchTST model achieving the second lowest average MAE among all the models.
Based on the average results across all stations, Informer with our proposed BFPF achieves the lowest MSE, indicating its superior overall predictive accuracy, while xLSTM ranks second. In terms of average MAE, TimeFilter attains the best performance, followed by PatchTST, demonstrating their strong capability in reducing absolute prediction errors.
V-D RQ2: Analysis from the Dataset Perspective
From a dataset perspective, the predictive performance across different sites reflects the distinct characteristics of their rainfall time series. J340 exhibits generally low errors, with MSE ranging from 1.30 to 1.50 and MAE from 0.30 to 0.50, indicating a relatively stable series with few extreme events, high data quality, and low prediction difficulty. ZIMM shows moderately high errors, with MSE exceeding 2.00, primarily due to pronounced seasonality and moderate precipitation events, reflecting a certain level of data complexity. P095 has low MSE values (0.3–0.5), suggesting sparse rainfall and relatively simple prediction conditions. MTLA demonstrates moderate errors, consistent with tropical rainfall concentrated in the wet season and minimal zero-inflation, making predictions relatively manageable. ARTA exhibits high errors, with model MSE generally above 4, indicating a challenging prediction task. BFTA shows relatively low MSE values between 0.75 and 0.90, reflecting localized rainfall patterns and moderate prediction difficulty. Finally, FLM5 presents near-zero errors (MSE $\approx$ 0.0000), as Antarctic precipitation is minimal and the time series is almost entirely zero, rendering the prediction task trivial.
V-E RQ3: Effect Analysis of Bi-Focus Precipitation Forecaster
TABLE III: Ablation analysis of the Non-Zero & Temporal Focus modules at the BFTA station. “24(2)” denotes an input sequence length of 24 and an output sequence length of 2; other notations follow the same convention. The red indicates the best-performing model.
Module 24(2) 24(4) 24(6)
Non-zero Focus Temporal Focus MSE MAE MSE MAE MSE MAE
0.6343 0.2004 0.8047 0.2358 0.7757 0.2470
0.3890 0.1615 0.4771 0.1844 0.4801 0.2016
0.3910 0.1655 0.4618 0.1784 0.4800 0.1964
0.3936 0.1592 0.4564 0.1800 0.4860 0.1929
TABLE IV: Ablation analysis of the Non-Zero & Temporal Focus modules at the P095 station. “24(2)” denotes an input sequence length of 24 and an output sequence length of 2; other notations follow the same convention. The red indicates the best-performing model.
Module 24(2) 24(4) 24(6)
Non-zero Focus Temporal Focus MSE MAE MSE MAE MSE MAE
0.3449 0.1031 0.3677 0.1015 0.3913 0.1287
0.0815 0.04971 0.0867 0.0537 0.0887 0.0522
0.0826 0.0510 0.0905 0.0537 0.0884 0.0571
0.0798 0.0467 0.0857 0.0584 0.0901 0.0565
Transformer-based models have gained widespread attention in recent years as powerful tools for sequential modeling across various domains. However, from Table II, it is evident that Transformer-based models exhibit relatively unstable performance in rainfall time-series forecasting. This limitation arises because standard Transformer architectures fail to explicitly capture key rainfall data characteristics, including high sparsity and rapid temporal decay.
To validate the effectiveness of the proposed BFPF module, we select Informer, one of the strongest Transformer-based models, as our baseline. To further assess the generalization capability of the module, ablation experiments are conducted on two representative stations, P095 and BFTA, under multiple forecasting horizons.